QUESTION:
A rectangle has a perimeter P of 58 inches. The length l is 1 inch more than 3 times the width w. Write and solve a system of linear equations to find the length and width of the rectangle.
WHAT I HAVE TRIED:
l=3w+1
58/2=29
29=???

Answer by richwmiller(17219)

l=3w+1 is good
but the
perimeter is 2w+2l=58
2w+2(3w+1)=58
2w+6w+2=58
8w=56
w=7
l=22
from your 29=??
2w+2l=58
w+l=29
29=3w+1+w
28=4w
7=w
l=22